OpenStudy (anonymous):

1. What are the next three terms in the sequence? -6, 5, 16, ,27 A - 38, 49, 60 B - 37, 47, 57 C - 36, 45, 54 D- 36, 46, 57

OpenStudy (anonymous):

common difference=5-(-6)=(16-5)=(27-16) can you find the next term and so on?

OpenStudy (anonymous):

you see every number is 11 more than the previous number so if number is 27 then what is the next number?

OpenStudy (anonymous):

I suck at math. Lol. @surjithayer

OpenStudy (anonymous):

But the answer would be 38. @surjithayer

OpenStudy (anonymous):

27+11=38 38+11=49 49+11=60